# Sensory Data Sharing in ISAC System: Fundamental Limits and Waveform Design

### v2

@@ 6       Step 4: Find $\lambda^{\star}$ by bisection searching ;
$\bullet$ Source distribution $P_{\mathsf{\tilde{S}}}(\tilde{s})$.
@@
Subsequently, given a specific channel input distribution $P_{\mathsf{X}}^{(\mu)}(x)$, our objective is to compute the estimate distribution $P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})$, which is essential for determining the rate-distortion function. However, deriving the explicit expression of the estimate distribution for an arbitrary state distribution $P_{\mathsf{S}}(s)$ is a challenging task. Instead, we can compute the estimate $\tilde{\mathsf{S}}$ and its distribution $P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})$ using formulas (4) and (11), respectively, by performing a sufficiently large number of random trials. For analytical convenience, we consider a Gaussian state distribution and a linear sensing model, which allows for explicit expressions to be derived.
Assume that the target’s state follows Gaussian distribution with $\mathcal{CN}(0,\nu_{s}^{2})$. Let us consider the following linear sensing model
@@
where $\mathsf{N}\sim\mathcal{CN}(0,1)$ represents Gaussian channel noise. In such a Gaussian linear model, the estimator (4) is indeed an MMSE estimator for each channel realization $x$. Therefore, the estimate $\mathsf{\tilde{S}}$ and the corresponding conditional probability distribution of $P_{\mathsf{\tilde{S}|X}}(\tilde{s}|x)$ can be expressed by
@@
Moreover, sensing cost is the MSE with the given realization $x$, which can be written by
@@
In this sub-problem, we aim to evaluate the minimum communication distortion $D_{c}^{(\mu)}$ achieved under the source distribution $P_{\mathsf{\tilde{S}}}^{(\mu)}(\tilde{s})$ and the channel capacity $I^{(\mu)}(\mathsf{X};\mathsf{Y})$. This scenario corresponds to a typical rate-distortion problem in lossy data transmission. The distortion-rate curve can be obtained by the optimization problem [29, Chapter 9]
@@
The unknown nuisance term $P_{\mathsf{\hat{S}}}(\hat{s})$ can be updated by
$P_{\mathsf{\hat{S}}}(\hat{s})=\sum_{\tilde{s}}P_{\mathsf{\tilde{S}}}(\tilde{s})Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s}).$ (43)
By setting an initial distribution $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$, the optimal solution of (41) can be obtained by updating (43) and (42) iteratively. The detailed derivations can be found in [29].
For a given slop $\lambda_{s}$, the tangent point $\big{(}D^{(\lambda_{s})}_{c},I^{(\lambda_{s})}(\mathsf{\tilde{S};\hat{S}})\big{)}$ of the distortion-rate curve can be obtained by solving problem (41). We may collect a series of distortion-rate data points to fit a distortion-rate function by varying the slop $\lambda_{s}$. Thus, the communication distortion constrained by a given channel capacity is attained by
@@
$\bullet$ CAS distortion.
In the outer loop, we may collect a set $\mathcal{D}$ of the CAS distortion $D^{(\mu)}=D_{s}^{(u)}+D_{c}^{(u)}$ and the associated channel input distribution $P_{\mathsf{X}}^{(\mu)}(x)$ by repeating the above steps with various penalty factor $\mu$. Therefore, the optimal channel distribution that minimizes the CAS distortion can be found by
[…]
@@
Input: $P_{\mathsf{\tilde{S}}}(\tilde{s})$, $I^{(u)}(\mathsf{X};\mathsf{Y})$, $K$, a set of the slops $\mathcal{A}$.
1 for $\lambda_{s}\in\mathcal{A}$ do
2       Initialize: Set $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$ as uniform distribution ;
3       for $k\leq K$ do
4             Step 1: Update output distritbution in (43) ;
5             Step 2: Update $Q_{\mathsf{\hat{S}|\tilde{S}}}(\hat{s}|\tilde{s})$ in (42) ;
6
7       end for
8      Step 3: Collect data $\mathcal{I}\leftarrow\big{(}D_{c}^{(\lambda_{s})},I^{(\lambda_{s})}(\mathsf{\tilde{S};\hat{S}})\big{)}$ ;
9
10 end for
11Step 4: Characterize the $D(R)$ function base on $\mathcal{I}$ ;
Output: Communication distortion in (44).
Algorithm 2 Communication Distortion Algorithm
@@ 4             Step 1: Update output distritbution in (43) ;
3       Step 2: Calculate estimate distribution $P^{(\mu)}_{\mathsf{\tilde{S}}}(\tilde{s})$ ;
@@ V Waveform Design for Gaussian Channel Input
The previous section illustrates that using Gaussian channel input may not be optimal for the CAS systems. Despite this limitation, the assumption of Gaussian signals is prevalent in communications research literature. In this section, we introduce an ISAC waveform design scheme specifically developed for multi-input multi-output (MIMO) systems, while adhering to the framework of Gaussian signaling.
@@ V-A System Model
with $\mathbf{z}=\text{vec}(\mathbf{Z}^{H})$, and $\mathbf{n}_{s}=\text{vec}(\mathbf{N}_{s}^{H})$. The BS estimates TRM $\tilde{\mathbf{s}}$ from the observations $\mathbf{z}$, then transmits it to the user through communication channel $\mathbf{H}_{c}$. Subsequently, the user can reconstruct $\hat{\mathbf{s}}$ from the communication received data $\mathbf{Y}$. Here, we adopt the quadratic distortion metric (MSE), which meets the separability condition as shown in (14). Before proceeding the ISAC waveform design, we make the following assumptions.
@@
1) Sensing distortion: For the Gaussian linear model (47), the optimal estimator (4), i.e., the well-known MMSE estimator, is leveraged to obtain the estimate $\tilde{\mathbf{s}}$ by
@@
2) Communication distortion: The estimate $\tilde{\mathbf{s}}$ in (48) follows the complex Gaussian distribution with $\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{s}}\otimes\mathbf{R}_{\tilde{s}})$, with the covariance matrix of [33]
@@
Let us temporarily omit the Kronecker product. For a Gaussian source $\mathcal{CN}(\mathbf{0},\mathbf{R}_{\tilde{s}})$, the rate-distortion function can be characterized by the following optimization problem [34]
@@
where $\mathbb{S}_{N_{t}}^{+}$ denotes the set of all $N_{t}\times N_{t}$ positive definite matrices. More details on (51) can be found in the proof of [34, Theorem 2].
@@
$C(\mathbf{R}_{x})=I(\mathbf{Y};\mathbf{X}|\mathbf{H}_{c})=\log\det\left(\frac{1}{\sigma^{2}_{c}}\mathbf{H}_{c}\mathbf{R}_{x}\mathbf{H}_{c}^{H}+\mathbf{I}_{M_{c}}\right).$ (52)
@@ V-B ISAC Waveform Design
Observe that the expressions (49) to (52) are all related to the covariance matrix $\mathbf{R}_{x}$ rather than the specific waveform matrix $\mathbf{X}$. Consequently, our ISAC waveform design problem can be transformed into determining the optimal covariance matrix that minimizes the CAS distortion. Regarding the system resource constraint, we only consider the transmit power at the BS. By substituting the expressions into the original problem $\mathcal{P}_{1}$ and introducing an auxiliary variable of positive definite matrix $\mathbf{D}$, the ISAC waveform design problem can be reformulated by
@@
where $P_{T}$ is the power budget. $\mathcal{P}_{2}$ is non-convex due to the nonlinear intermediate variables $\log\det(\mathbf{R}_{\tilde{s}})$ and $\mathbf{R}_{\tilde{s}}$. To relax $\mathcal{P}_{2}$ into a convex problem, we employ an SCA technique based on Taylor series expansion. To be specified, for any given point $\mathbf{R}_{0}$, the matrix $\mathbf{R}_{\tilde{s}}$ can be approximated by
@@ VI Simulation Results
The simulation results are divided into two sections. The first section evaluates the optimal distribution for the ISAC channel input in scalar cases, demonstrating the effectiveness of the proposed MSST and the two-step BA-based optimal search algorithm. Moreover, the unique distortion-capacity (D-C) curves of the CAS system are provided to illustrate the performance tradeoff between S&C processes. The second section presents the results of the optimal waveform design for the Gaussian MIMO channel, showing its superiority over the S&C-optimal schemes.
@@ VI-A Optimal Distribution for ISAC Channel Input
Figure 4: The CAS distortion versus the communication channel capacity.
Figure 5: The channel input distribution of S&C-optimal and CAS-optimal.
VI-A1 The D-C curves and the optimal distribution of ISAC channel input
Fig. 3 presents the D-C curves for both S&C processes. As expected, sensing distortion $D_{s}$ increases as the values of $\mu$ decrease, implying that the estimation performance at the BS becomes less significant. Conversely, the communication channel capacity increases accordingly, leading to a reduction in communication distortion $D_{c}$. Consequently, these opposing trends in S&C distortions result in an uncertain CAS distortion. In Fig. 4, we show the resultant D-C curve for the CAS systems with different sensing SNRs. A CAS-optimal point that minimizes the CAS distortion is evident, highlighting a unique phenomenon in CAS systems compared to the existing studies [14, 15].
Fig. 5 depicts the specific channel input distributions corresponding to three special points in Fig.4, namely, the S&C-optimal and CAS-optimal points. The channel capacity is maximized without the sensing constraint when $\mu=0$. It is well-known that Gaussian channel input achieves the maximum MI for the Gaussian channels. By contrast, at the sensing-optimal point, as $\mu$ becomes sufficiently large, the channel input exhibits a 2-ary pulse amplitude modulation. These results are consistent with the findings in [12]. However, the unique CAS-optimal distribution, presenting a compromise between the S&C-optimal distributions, is attained at $\mu=2.5$.
@@ VI-A1 The D-C curves and the optimal distribution of ISAC channel input
Figure 7: The CAS-optimal distribution for various state variances.
@@ VI-A2 The impact of the target’s prior variance
Subsequently, we investigate the impact of the target’s state variance by selecting the values of $\nu_{s}^{2}\in\{0.1,0.2,0.5,5\}$. Due to the significant differences in the order of magnitude of the values for MI and sensing distortions in (32), we adjust the interval of $\mu$ for each state variance rather than using a fixed interval. The essential purpose of the CAS systems is to reduce the uncertainty of the target’s prior information, specifically the variance of the state’s prior distribution.
In Fig. 6, each point represents a lower CAS distortion compared to the prior variance, even at the S&C-optimal points. This reduction is attributed to the resource multiplexing gain achieved through the dual-functional signaling strategy. Additionally, we observe that the CAS-optimal point gradually shifts from sensing-optimal to communication-optimal as variance increases. A small variance indicates relatively accurate prior information about the target’s state, necessitating an ISAC channel input with strong sensing capabilities for better accuracy performance. Conversely, as variance increases, greater emphasis on communication performance is required to convey more information.
We provide the detailed distributions of the CAS-optimal channel input for different state variances. For a small variance (e.g., $\nu_{s}^{2}=0.1$), the sensing distortion becomes small accordingly. It is at least less than the prior variance; otherwise, the user can ‘guess’ the state information in terms of the prior distribution. Therefore, a large weighted factor is required, e.g., $\mu=690$, to balance the values of MI and sensing distortion in (32). In this case, the CAS-optimal distribution aligns with the sensing-optimal distribution, specifically a standard 2-ary pulse amplitude modulation. As anticipated, the optimal $\mu$ values decrease with increasing state variance, and the CAS-optimal distribution changes accordingly.
@@ VI-B1 Waveform design for a 2-D special case
We begin with a two-dimensional (2-D) special case where the state covariance matrix, $\bm{\Sigma_{s}}=\text{diag}\{a_{1},a_{2}\}$, is assumed to be diagonal. In this scenario, the global optimal solution of problem $\mathcal{P}_{2}$ can be attained through a 2-D exhaustive search, as outlined in [11]. Thus, this optimal solution  [^4] The optimal eigenspace of $\mathbf{R}_{x}$ must align with the communication eigenspace of $\mathbf{H}_{c}^{H}\mathbf{H}_{c}$ as the sensing eigenspace is an arbitrary unitary matrix. Thus, the original problem can be reduced to finding two optimal eigenvalues over 2-D grids, with the sum of each 2-D point equal to $P_{T}$. can serve as a benchmark for comparing the superiority of our method. Accordingly, the numbers of the Tx and S&C Rx antennas are set by $N_{t}=M_{c}=M_{s}=2$.
Fig. 8 illustrates the impact of sensing channel quality on the achievable average distortion. The sensing SNRs vary from 00 dB to $20$ dB with a spcing $2$ dB. Besides the global optimal benchmark, the S&C optimal waveform schemes refer to the solutions that minimize the MMSE in (49) and maximize the channel capacity in (52), respectively. At a glance, we observe that the achievable average distortion decreases with increasing sensing channel SNRs. Furthermore, the sensing-optimal scheme is preferable for acquiring accurate target information when the communication channel is sufficiently good at $\text{SNR}_{c}=5$ dB. However, the sensing-optimal scheme may cause significant performance degradation as the communication SNR decreases. Particularly, the sensing-optimal curve even tends to be flat at $\text{SNR}_{c}=-5$ dB, implying that a poor communication channel severely limits performance improvement. This clearly highlights the unique tadeoff between S&C processes in CAS systems. Finally, it is worth noting that the proposed algorithm achieves satisfactory performance compared to the global optimal solution.
Figure 9: The impact of the state variances on the CAS performance, $\text{SNR}_{s}=5$ dB, $N_{t}=M_{c}=M_{s}=2$.

### v3

@@ 6 Step 4: Find $\lambda^{\star}$ by bisection searching ;
[…]
@@
[…]
Assume that the target’s state follows Gaussian distribution with $P_{\mathsf{S}}(s)=\mathcal{CN}(0,\nu_{s}^{2})$. Let us consider the following linear sensing model
@@
where $\mathsf{N}\sim\mathcal{CN}(0,1)$ represents Gaussian channel noise. In such model, the estimator (4) is indeed an MMSE estimator for each channel realization $x$. Therefore, the estimate $\mathsf{\tilde{S}}$ and the corresponding conditional probability distribution of $P_{\mathsf{\tilde{S}|X}}(\tilde{s}|x)$ can be expressed by
@@
Moreover, sensing cost $e(x)$ is MSE with given realization $x$, which can be written by
@@
[…]
1 for $\lambda_{s}\in\mathcal{A}$ do
[…]
3 for $k\leq K$ do
4    Step 1: Update output distritbution in (42) ;
[…]
6
7   end for
8 Step 3: Collect data $)\mathcal{I}\leftarrow\big(D_{c}^{(\lambda_{s})},I^{(\lambda_{s})}(\mathsf{\tilde{S};\hat{S}}$$\big)$ ;
9
10 end for
11Step 4: Construct $D(R)$ function by data fit in $\mathcal{I}$ ;
Output: Communication distortion in (43).
Algorithm 2 Communication Distortion Algorithm
@@ 4    Step 1: Update output distritbution in (42) ;
[…]
@@
[…]
For a given slop $\lambda_{s}$, the tangent point $)\big(D^{(\lambda_{s})}_{c},I^{(\lambda_{s})}(\mathsf{\tilde{S};\hat{S}}$$\big)$ of the distortion-rate curve can be obtained by solving problem (40). We may collect a series of distortion-rate data points to fit a distortion-rate function by varying the slop $\lambda_{s}$. Thus, the communication distortion constrained by a given channel capacity is attained by
@@
$\bullet$ SDAS distortion.
In the outer loop, we may collect a set $\mathcal{D}$ of the SDAS distortion $D^{(\mu)}=D_{s}^{(u)}+D_{c}^{(u)}$ and the associated channel input distribution $P_{\mathsf{X}}^{(\mu)}(x)$ by repeating the above steps with various penalty factor $\mu$. Therefore, the optimal channel distribution that minimizes the SDAS distortion can be found by
$D^{(\mu^{\star})}=\mathop{\arg\min}\limits_{D\in\mathcal{D}}\kern 2.0ptD,\kern 5.0ptP_{\mathsf{X}}^{(\star)}(x)=P_{\mathsf{X}}^{(\mu^{\star})}(x).$ (44)
@@
[…]
@@ V Waveform Design for Gaussian Channel Input
The previous section illustrates that using Gaussian channel input may not be optimal for the SDAS systems. Despite this limitation, the assumption of Gaussian signals is prevalent in communications research literature. In this section, we introduce an ISAC waveform design scheme specifically developed for multi-input multi-output (MIMO) systems, while adhering to the framework of Gaussian signaling. To proceed, we need to derive the explicit expressions of the variables in $\mathcal{P}_{0}$ for the MIMO case.
@@ V-A System Model
with $\mathbf{z}=\text{vec}(\mathbf{Z}^{H})$, and $\mathbf{n}_{s}=\text{vec}(\mathbf{N}_{s}^{H})$. The BS yields an estimate of the TRM, namely, $\tilde{\mathbf{s}}$, from the observations $\mathbf{z}$, then transmits it to the user through communication channel $\mathbf{H}_{c}$. Subsequently, the user can reconstruct $\hat{\mathbf{s}}$ from the communication received data $\mathbf{Y}$. Here, we adopt the quadratic distortion metric (i.e., MSE), which meets the separable condition in (13). Before proceeding the ISAC waveform design, we make the following assumptions.
@@
1) Sensing distortion $D_{s}$: For the Gaussian linear model (46), the optimal estimator (4), i.e., the well-known MMSE estimator, is leveraged to estimate $\tilde{\mathbf{s}}$ by
@@
2) Rate distortion function $R(D_{c})$: The estimate $\tilde{\mathbf{s}}$ in (47) follows the complex Gaussian distribution $\mathcal{CN}(\mathbf{0},\mathbf{I}_{M_{s}}\otimes\mathbf{R}_{\tilde{s}}), with$ the covariance matrix of [37]
@@
[…]
@@
where $\mathbb{S}_{N_{t}}^{+}$ denotes the set of all $N_{t}\times N_{t}$ positive definite matrices. Denote $\mathbf{D}^{\star}$ as the solution of (50), the communication distortion $D_{c}$ and the required rate $R(D_{c})$ can be obtained by
$D_{c}=M_{s}\text{Tr}\left(\mathbf{D}^{\star}\right),\kern 5.0ptR(D_{c})=M_{s}\log\left(\frac{\det(\mathbf{R}_{\tilde{s}})}{\det(\mathbf{D}^{\star})}\right),$ (51)
where $M_{s}$ is due to the Kronecker product. The derivations of (50) can be found in the proof of [38, Theorem 2]  [^4] Our result is a straightforward corollary with the semantic distortion equal to zero in Theorem 2..
@@
$C(\mathbf{R}_{x})=\log\det\left(\frac{1}{\sigma^{2}_{c}}\mathbf{H}_{c}\mathbf{R}_{x}\mathbf{H}_{c}^{H}+\mathbf{I}_{M_{c}}\right).$ (52)
@@ V-B ISAC Waveform Design
Observe that the expressions (48) to (52) are all related to the covariance matrix $\mathbf{R}_{x}$ rather than the specific waveform matrix $\mathbf{X}$. Consequently, our ISAC waveform design problem can be transformed into determining the optimal covariance matrix that minimizes the SDAS distortion. Regarding the system resource constraint, we only consider the transmit power at the BS. By substituting the expressions into the original problem $\mathcal{P}_{1}$ and introducing an auxiliary variable of positive definite matrix $\mathbf{D}$, the ISAC waveform design problem can be reformulated by
@@
[…]
@@ VI Simulation Results
The simulation results are divided into two sections. The first section evaluates the optimal distribution for the ISAC channel input in scalar cases, demonstrating the effectiveness of the proposed MSST and the two-step BA-based optimal search algorithm. Moreover, the unique distortion-capacity (D-C) curves of the SDAS system are provided to illustrate the performance tradeoff between S&C processes. The second section presents the results of the optimal waveform design for the Gaussian MIMO channel, showing its superiority over the S&C-optimal schemes.
@@ VI-A Optimal Distribution for ISAC Channel Input
Figure 4: The SAS distortion versus the communication channel capacity.
Figure 5: The channel input distribution of S&C-optimal and SAS-optimal.
VI-A1 The Distortion-Capacity (D-C) curves and the optimal distribution of ISAC channel input
Fig. 3 presents the D-C curves for both S&C processes. As expected, sensing distortion $D_{s}$ increases as the values of $\mu$ decrease, implying that the estimation performance at the BS becomes less significant. Conversely, the communication channel capacity increases accordingly, leading to a reduction in communication distortion $D_{c}$. Consequently, these opposing trends in S&C distortions result in an uncertain SDAS distortion. In Fig. 4, we show the resultant D-C curve for the SDAS systems with different sensing SNRs. A SDAS-optimal point that minimizes the SDAS distortion is evident, highlighting a unique phenomenon in SDAS systems compared to the existing studies [20, 22].
Fig. 5 depicts the specific channel input distributions corresponding to three special points in Fig. 4, namely, the S&C-optimal and SDAS-optimal points. The channel capacity is maximized without the sensing constraint when $\mu=0$. It is well-known that Gaussian channel input achieves the maximum MI for the Gaussian channels. By contrast, at the sensing-optimal point, as $\mu$ becomes sufficiently large, the channel input exhibits a 2-ary pulse amplitude modulation. These results are consistent with the findings in [19]. However, the unique SDAS-optimal distribution, presenting a compromise between the S&C-optimal distributions, is attained at $\mu=2.5$.
@@ VI-A1 The Distortion-Capacity (D-C) curves and the optimal distribution of ISAC channel input
Figure 7: The SDAS-optimal distribution for various state variances.
@@ VI-A2 The impact of the target’s prior variance
Subsequently, we investigate the impact of the target’s state variance by selecting the values of $\nu_{s}^{2}\in\{0.1,0.2,0.5,5\}$. Due to the significant differences in the order of magnitude of the values for MI and sensing distortions in (31), we adjust the interval of $\mu$ for each state variance rather than using a fixed interval. The essential purpose of the SDAS systems is to reduce the uncertainty of the target’s prior information, specifically the variance of the state’s prior distribution.
In Fig. 6, each point represents a lower SDAS distortion compared to the prior variance, even at the S&C-optimal points. This reduction is attributed to the resource multiplexing gain achieved through the dual-functional signaling strategy. Additionally, we observe that the SDAS-optimal point gradually shifts from sensing-optimal to communication-optimal as variance increases. A small variance indicates relatively accurate prior information about the target’s state, necessitating an ISAC channel input with strong sensing capabilities for better accuracy performance. Conversely, as variance increases, greater emphasis on communication performance is required to convey more information.
We provide the detailed distributions of the SDAS-optimal channel input for different state variances. For a small variance (e.g., $\nu_{s}^{2}=0.1$), the sensing distortion becomes small accordingly. It is at least less than the prior variance; otherwise, the user can ‘guess’ the state information in terms of the prior distribution. Therefore, a large weighted factor is required, e.g., $\mu=690$, to balance the values of MI and sensing distortion in (31). In this case, the SDAS-optimal distribution aligns with the sensing-optimal distribution, specifically a standard 2-ary pulse amplitude modulation. As anticipated, the optimal $\mu$ values decrease with increasing state variance, and the SDAS-optimal distribution changes accordingly.
@@ VI-B1 Waveform design for a 2-D special case
We begin with a two-dimensional (2-D) special case where the state covariance matrix, $\bm{\Sigma_{s}}=\text{diag}\{a_{1},a_{2}\}$, is assumed to be diagonal. In this scenario, the global optimal solution of problem $\mathcal{P}_{2}$ can be attained through a 2-D exhaustive search, as outlined in [18]. Thus, this brute force solution  [^5] The optimal eigenspace of $\mathbf{R}_{x}$ must align with the communication eigenspace of $\mathbf{H}_{c}^{H}\mathbf{H}_{c}$ as the sensing eigenspace is an arbitrary unitary matrix. Thus, the original problem can be reduced to finding two optimal eigenvalues over 2-D grids, with the sum of each 2-D point equal to $P_{T}$. can serve as a benchmark for comparing the superiority of our method. Accordingly, the numbers of the Tx and S&C Rx antennas are set by $N_{t}=M_{c}=M_{s}=2$.
Fig. 8 illustrates the impact of sensing channel quality on the achievable average distortion. The sensing SNRs vary from 0 dB to $20$ dB with a spcing $2$ dB. Besides the global optimal benchmark, the S&C optimal waveform schemes refer to the solutions that minimize the MMSE in (48) and maximize the channel capacity in (52), respectively. At a glance, we observe that the achievable average distortion decreases with increasing sensing channel SNRs. Furthermore, the sensing-optimal scheme is preferable for acquiring accurate target information when the communication channel is sufficiently good at $\text{SNR}_{c}=5$ dB. However, the sensing-optimal scheme may cause significant performance degradation as the communication SNR decreases. Particularly, the sensing-optimal curve even tends to be flat at $\text{SNR}_{c}=-5$ dB, implying that a poor communication channel severely limits performance improvement. This clearly highlights the unique tadeoff between S&C processes in SDAS systems. Finally, it is worth noting that the proposed algorithm achieves satisfactory performance compared to the global optimal solution.
Figure 9: The impact of the state variances on the SAS performance, $\text{SNR}_{s}=5$ dB, $N_{t}=M_{c}=M_{s}=2$.